Constants Part A A wire 6.20 m long with diameter of 1.85 mm has a resistance of 0.0310 Ω Find the resistivity of the material of the wire. ΑΣ AMD 1.3.10- 2-m Submit Previous Answers Request Answer X Incorrect; Try Again; 5 attempts remaining
Answer #1

Screenshot_20210517_212250.JPG

answered by: ANURANJAN SARSAM
Answer #2

here given data,  

resistance, R =0.0310 Ω

length lo wire, L= 6.20 m

diameter of wire , D = 1.85 mm

= 0.00185 m [ here 1 meter = 1000 mm,1.85/1000 = 0.00185 m ]

here area of wire ,

A = π D2 / 4

= 3.1428 * ( 0.00185 )2 /4

= 3.1428 * 0.0000034225 / 4

= 0.00001075623 / 4

A = 0.00000268905 m2

here , we have formula for resistivity,

  ρ = RA/ L

= 0.0310 *  0.00000268905 / 6.20

= 0.0000001060975 / 6.20

ρ  = 0.0000000171125 Ω m

   ρ = 1.71125 × 10-8   Ω m

Here resistivity of wire is  ρ  = 1.71125 × 10-8   Ω m